There are two tests for a particular antibody. Test A gives a correct result 95% of the time. Test B is accurate 89% if the time. If a patient is given both tests, find the probability that:
At least one of the tests gives the correct result.

Answer by ewatrrr(24785)

pA(correct) = .95, pB(correct) = .89
patient is given both tests
P(at least one correct) = 1 - P(neither correct) = 1 - (.05)(.11) = 1 - .0055 = . 9945 0r 99.45%